Penguin Ice Inc. was formed on June 30, 2015, through the merger of Penguin Corp. and Ice Inc. Penguin Ice issued a total of 2,500,000 shares of common stock to owners of the merged companies. The new company had the following transactions during 2017. On April 1, 2017, the company issued an additional 500,000 shares of stock for cash. On July 1, 2017, Penguin Ice issued $1 million of 15-year, 6% convertible bonds at par. Each $1,000 bond converts to 25 shares of common at any interest date. None of the bonds have been converted to date. Penguin Ice is preparing its annual report for the fiscal year ending December 31, 2017, and will report after-tax net income of $13,600,000. The tax rate is 40%.

Instructions Determine the following for 2017. (a) The number of shares to be used for calculating: (1) Basic earnings per share. (2) Diluted earnings per share (b) The earnings figures to be used for calculating: (1) Basic earnings per share. (2) Diluted earnings per share
Answer #1

Please find below table useful to compute desired results: -

А B 1 After tax net Income 13600000 2 Total outstanding shares =2500000+500000 3 Basic EPS =B1/B2 4 5 No. of Shares on Conver

End results would be as follows: -

B 13,600,000 3,000,000 4.53 A 1 After tax net Income 2 Total outstanding shares 3 Basic EPS 4 5 No. of Shares on Conversion 6